OpenStudy (anonymous):

What is the 11th term of the following geometric sequence? 7/9, -7/3, 7, -21, 63..

OpenStudy (anonymous):

45,927

OpenStudy (mathstudent55):

\(a_1 = a_1\) \(a_2 = a_1 \times r\) \(a_3 = a_2 \times r = a_1 \times r^2\) \(a_4 = a_3 \times r = a_1 \times r^3\) \(...\) \(a_n = a_1 r^{n - 1} \) \(r = -\dfrac{7}{3} \div \dfrac{7}{9} = -\dfrac{7}{3} \times \dfrac{9}{7} = -3 \) \(\large a_n = \dfrac{7}{9} \times (-3)^{10 - 1} \)

OpenStudy (mathstudent55):

Last line above should read: \(\large a_{11} = \dfrac{7}{9} \times (-3)^{11 - 1} \) Now evaluate this expression to find \(a_{11} \).